OpenStudy (anonymous):

what would be the order pair if you was to sketch a graph of y=f(x) when f(x) = -4x+2

OpenStudy (zzr0ck3r):

it asks for "the" ordered pair?

OpenStudy (anonymous):

it is asking me to sketch a graph of y=f(x). f(x)= -4x+2

OpenStudy (zzr0ck3r):

what Is the slop?

OpenStudy (zzr0ck3r):

slope*

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

Its a equation of straight line . Now can u proceed ??

OpenStudy (anonymous):

I have not a single clue on how to do this one

OpenStudy (zzr0ck3r):

then make a table and pick some x values, plug them in and get the y values. like x = 1 then y = -2 and x=2 then y = -6 now plot the points (1,-2) and (2,-6) and draw a straight line through them

OpenStudy (anonymous):

In General Any equation in form of y=mx+c is a straight line with m= slope & c= intercept on y axis here m=-4& c=2 Now to plot any graph Put x=0 & then find y=f(x) here Then x,y) u just found will be an ordered pair Similarly y=0 & then find x here This (x,y) u just found will be another ordered pair